how do i find the slope -intercept form of the equation of the line that passes through(5,-1)(-5,5)
Answer by stanbon(57984)

find the slope -intercept form of the equation of the line that passes through(5,-1)(-5,5)
----------------------
slope = (change in y)/(change in x) = (5--1)/(-5-5) = 6/-10 = -(3/5)
----------------------------------------
Equation Form: y = mx+b
Solve for "b"
5 = (-3/5)(-5)+b
5 = 3+b
b = 2
--------------------------
Equation:
y = (-3/5)x+2
==================
Cheers,
stan H.